What is the minimum value of the given function? 𝑓(𝑥)=6𝑥^2 −12𝑥+𝑘, 0≤𝑥≤10, 𝑘 is a constant.
(A)𝑓(0) (B) 𝑓(10) (C) 𝑓(0) + 𝑘 (D) 𝑓(1) + 𝑘 (E) none of the above.

Found 2 solutions by ewatrrr, Alan3354:
Answer by ewatrrr(24785)

𝑓(𝑥)=6𝑥^2 −12𝑥+𝑘, 0≤𝑥≤10, 𝑘 is a constant.
Find minimum value of f(x)
minimum at x = -b/2a = 1
𝑓(𝑥)=6𝑥^2 −12𝑥+𝑘
f(1) = 6 - 12 + k = k - 6

Answer by Alan3354(69443)